OpenStudy (anonymous):

What is the 56th term of the sequence below? -29,-29,-25,-21,-17......

OpenStudy (anonymous):

The formula for this is a_n=a_1+(n-1)d. So the numbers into that.

OpenStudy (anonymous):

\[a _{n}=a _{1}+(n-1)d\]

OpenStudy (anonymous):

im sorry im not good at this can u explain

OpenStudy (anonymous):

a_n is the term you are trying to find and a_1 is the first term. d is the difference between the numbers. Just to be sure, there are not two -29s at the beginning of the sequence right

OpenStudy (anonymous):

ok yes thank you

OpenStudy (anonymous):

this should make the equation \[a _{56}=-29+(n-1)(-4)\]

OpenStudy (anonymous):

Then, solve

OpenStudy (anonymous):

56?

OpenStudy (anonymous):

Sorry, i gave you the wrong formula. Sorry, it is really S_n = n/2 [2a_1 + (n – 1)d] The reason is because you dont know the last term

OpenStudy (anonymous):

Does it say how many terms are in the sequence?

OpenStudy (anonymous):

oh ok so N would be 56 right?

OpenStudy (anonymous):

A is -29 right?

OpenStudy (anonymous):

i dont know what d would be

OpenStudy (anonymous):

Never mind, I gave you the right equation first. n would be 56 because that could be the last term. It doesn't matter. I am sorry, I just learned this and I am trying to get better practice

OpenStudy (anonymous):

what would D be?

OpenStudy (anonymous):

d is 4 because -25-(-29) is -25+29 which is 4 a56=−29+(56−1)(4) a56=-29+220 a56=191

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

Sorry about the confusion :(

OpenStudy (anonymous):

its fine!! :)